In a class 40% of student like music 3/15 like dance 1/10 like cricket and not painting. The remaining 40 student like painting. Find the no of student who like dance.
1
Expert's answer
2011-06-29T09:23:29-0400
Let's denote the whole amount of students as X. Thus
40/100 X = 0.4X like musik, 3/15X = 0.2X& like dance, 1/10X = 0.1X & like cricket, and X(1 - 0.4 - 0.2 - 0.1) = 40& like painting.
Let's simplify the lat expression:
X(1 - 0.4 - 0.2 - 0.1) = 40
0.3 X = 40
X = 40/0.3 = 133.33
May be there is some mistake in the conditions of the task because the number& of students should be the whole number. The number of students who like dance is 0.2X.